OpenStudy (anonymous):

Find the derivative of h(x)=tan^3x

OpenStudy (jdoe0001):

use the chain rule

OpenStudy (anonymous):

\[\frac{d}{dx} (\tan(x))^3\] You have a case where: \[\frac{d}{dx}(f(g(x)) = f'(g(x))g'(x)\] , the chain rule, applies. Using f(x) = x^3 and g(x) = tan(x) see if you can figure it out.

OpenStudy (anonymous):

3tan^2x sec^2x

OpenStudy (jdoe0001):

yeap